OpenStudy (anonymous):

I HAVE A FINAL TOMORROW, PLEASE HELP ME Simplify: 7i^2-i^3

OpenStudy (swissgirl):

Is it i as in \( \sqrt{-1}\)?

OpenStudy (anonymous):

isn't I -1??????????????????

OpenStudy (swissgirl):

No \[i=\sqrt{-1}\]

OpenStudy (anonymous):

oh sorry LOL I thought you said i^2

OpenStudy (anonymous):

and yeah, you're right

OpenStudy (anonymous):

so, how do you do this?

OpenStudy (swissgirl):

ok i^2=-1 and \[i^3=\sqrt{i*i*i}=\sqrt{i^2*i}=-1\sqrt{-1}=-1i\]

OpenStudy (anonymous):

UHHHHHHHHHHHH the answer is 7+i

OpenStudy (swissgirl):

hmmmm 7i-i=7(-1)-(-1i)=??? Watch ur signs

OpenStudy (anonymous):

lol sorry the problem is 7i^4 not 7i^2

OpenStudy (anonymous):

sorry

OpenStudy (swissgirl):

Ok gotcha lol

OpenStudy (anonymous):

ok well i get how to get the 7 cause -1 * -1 is a positive, but I still don't get the the i is a positive as well you said it was -1i

OpenStudy (swissgirl):

nah cuz look 7(1)-(-1i)=7+i

OpenStudy (anonymous):

OHHHHHHHHHHHHHHHHHHHHHHH OK, I GET IT THANK-YOU SO MUCH

OpenStudy (unklerhaukus):

7i^2-i^3 factor = i^2(7-i) use i^2=-1 = -(7-i) = -7+i